# Extended-Use Designs on Very Large Online Platforms

### v1

@@ 2.3. Adolescent Development and Manipulative Design
This is true for users of all ages but particularly true for teenagers, as teens’ reduced impulse control and heightened sensitivity to rewards and social feedback make them particularly vulnerable to pressure and distraction (digital or otherwise).During adolescence, executive functions such as impulse inhibition, flexible problem-solving, and abstract reasoning develop at different rates between teens, with most of these higher-order thinking skills reaching maturity in the mid-20s. Therefore, teens are more susceptible to digital designs that exert manipulation through curiosity and novelty/risk-seeking, social pressure, and fear of missing out, with significant inter-individual variability  (of Sciences Engineering and Medicine, 2024). Teens’ accounts of their experiences with technology reflect frustration with attention-economy designs; more than half of teens say that they are online “almost constantly,” and approximately 40% of teens say they wish they spent less time using devices and social media (Center, 2018).
@@
As a result, many teens report attempting to cut back on these experiences. They delete apps that are too aggressive in demanding their attention (Landesman et al., 2024) and turn their account passwords over to their friends to lock them out of platforms that they find too addictive (Baumer et al., 2013). They add friction by burying apps in folders  (Böhmer and Krüger, 2013), setting do-not-disturb times  (Natarajan, 2024), and accessing platforms through browsers. But despite their makeshift strategies to resist the design tactics of the attention economy, teens still struggle to change their habits (Tran et al., 2019), suffer from device-based sleep disruption (Woods and Scott, 2016; Viner et al., 2019), and struggle with distraction from activities they value, like homework and time spent with friends and loved ones  (Yardi and Bruckman, 2011). Parents of younger children report similar experiences and say that they frequently negotiate or argue with their child about transitioning away from media, especially media that uses engagement-prolonging designs like autoplay (Davis et al., 2019; Hiniker et al., 2016).
@@
In the midst of the ongoing debate regarding social media, technology, and youth mental health, there has been increasing focus on the role of design features as a mechanism linking platform use to reduced wellbeing (Abuse and, SAMHSA). In particular, designs such as autoplay , endless scroll, and time pressure were described by the U.S. White House Interagency Kids Online Health and Safety Task Force as warranting modifications by platforms in order to reduce compulsive and excessive use of technology. More research is needed to describe extended use manipulative designs and their impact on children and teens.
@@ 3. Method
We conducted an analysis of the web and app interfaces of all 17 VLOPs identified by the Digital Services Act (DSA) (list available here), attending specifically to the presence of EUDs. We conducted this work in two phases (see Figure 2); in the first phase, we performed an inductive-deductive thematic analysis to identify and categorize the EUDs present in these interfaces, examining these designs in the context of the extant literature, and organizing them by the deeper strategies that underlie them. In the second phase, we conducted a structured content analysis of all VLOP interfaces to document all instances of the EUDs we identified in phase 1.
@@ 3.1. Sampling
We engaged extensively with each VLOP web interface ($N=17$) and app interface ($N=16$), in each case posing as a teenage user located in the European Union. To observe websites, we used the Chrome browser in Guest Mode on a Windows 11 laptop. To observe mobile apps, we used a factory-reset smartphone running Android 9, equipped with an active SIM card. These devices were selected because they run the most widely used operating systems in Europe (StatCounter, 2024a, b). We created two new Google accounts, which we used to access all VLOPs (we refer to these as the “primary account,” which was used to engage with all features, and the “secondary account,” which was used to enable the primary account to engage with social features that required a partner). The only exception to the above procedures was our analysis of the Apple App Store interface, which we observed using iOS devices and accounts due to the platform’s restrictions.
Following the method developed by Di Geronimo and colleagues (Di Geronimo et al., 2020), we collect screen-capture videos of each service on each modality that we could later examine in detail for evidence of EUDs. Both mobile and desktop interactions were recorded using Screen Recorder over Wi-Fi. Since platform companies may conduct A/B testing their products, the exact behaviors or interfaces we observed in a platform may not represent what all users would have encountered during the study period (the spring and summer of 2024).
@@
    Download App and Create an Account: We began by downloading the app and creating a new account with our teenager persona. The primary Google account was set up with the profile of a 17-year-old girl. Additionally, we registered a secondary account to interact with the primary account, as detailed in step four.
@@
While the screen recordings contained the visual captures of EUD and interactions thereof, the audio served as complementary material, which allowed the coders to get a detailed impression of reviewers’ perception and judgement of a scene. We captured both screenshot and video records of the the user interfaces during these exploration and complied annotated images in a FigJam  (Fig, 2012) collective virtual workspace as structured data samples.
3.2. Content Analysis: Identifying the EUDs used by VLOPs
We then conducted a two-part analysis to identify the EUDs used by VLOPs. We first performed an inductive-deductive thematic analysis to develop a taxonomy of the EUDs used by VLOPs. We then conducted a structured content analysis to apply it comprehensively across platforms. Here, we describe these stages in detail.
@@ 3.2. Content Analysis: Identifying the EUDs used by VLOPs
    Inductive-Deductive Coding. We began our exploration with guidance from existing literature on attention-capture. In particular, we used the categories described by Radesky et al. (Radesky et al., 2022), Monge Roferello et al. (Monge Roffarello et al., 2023), and Flayelle et al. (Flayelle et al., 2023) as sensitizing concepts (Tracy, 2013) during our review of VLOP interfaces. We also allowed for emergent themes in EUDs and noted designs that did not fit neatly into these pre-existing taxonomies. Four authors independently reviewed the VLOP interfaces using this inductive-deductive approach and documented potential EUDs in a collective spreadsheet while interacting with VLOPs as described above. We also shared and organized exemplar screenshots in a collaborative Figma (Fig, 2012) board. Consistent with prior work (Di Geronimo et al., 2020; Gunawan et al., 2021), we made note of potential EUDs regardless of our perceptions of the designer’s intent, allowing us to include EUDs that may not reflect malicious intentions. The team met regularly over several months, and with each iteration brought new examples to discuss. During discussions, the team compared and contrasted designs, discussed possible groupings, assessed design examples against existing taxonomies, and refined possible category definitions. This process 38 inductive and 85 deductive EUD codes and associated examples.
@@
    Hierarchical Affinity Diagramming. The research team then conducted more formal affinity diagramming (Blandford et al., 2016) to restructure the codes and create an initial codebook. We structured our initial codebook into three levels: the psychological mechanisms through which the EUD design operates (high level), the recurring EUD design tactics (medium level), and the specific EUD design elements or interface components (low level). We added definitions and descriptions for each item at each level.
@@
    Interrater Reliability Testing. Two authors then independently reviewed two randomly selected VLOPs and applied the full set of codes across the interface. For each EUD, reviewers were required to note the code number of the mechanism and pattern, as well as specify the features exhibiting the EUD (e.g., ”1(a) gifting option”). The coding criteria focused on the distribution of EUDs across screenshots or interactions extracted from the recordings, rather than the frequency of EUD occurrences. In those cases where an aspect of the interface was deemed a potential EUD but was not present within the existing codebook, reviewers proposed and added a new code as a comment with a visual example. Any confusion about specific cases or the taxonomy itself was recorded as a comment. Inter-rater reliability (IRR) was established and measured following Lombard et al.  (Lombard et al., 2002). The results of the first round yielded an IRR score below 0.6, indicating that the initial codebook (v1) fell below the minimum validity standard. Subsequently, the research team collaboratively compared the two internal coding results, discussed points of disagreement, unresolved questions, and confusion comments. This led to the development of codebook (v2), which included a refined definitions and names for each code, and the modification of codes. Codebook(v2) includes 71 EUD features in total. The same two researchers conducted a second round of internal testing with two different randomly chosen platforms, adhering to the same tasks with codebook (v2). The IRR score for this round was 0.715, exceeding the standard of 0.6 and indicating that the revised coding scheme and codebook (v2) were clear and interpretable enough to be the final version.
@@
    Assessing the Prevalence of EUDs across VLOPs. Finally, two authors independently coded all recordings from all VLOP interfaces for each of the EUDs we documented. They then compared their coding results and came up with the final data.
Figure 3. Final version of EUD codebook, more details can be found here.
4. Results Part 1: The EUDs Used by VLOPs and the Underlying Strategies Behind Them
We found that the EUDs that VLOPs employ fall into four broad categories, defined by the underlying manipulation strategy the design draws upon. These include: 1) artificial obligations manufactured by the platform that pressure users into engaging, 2) reward- or pleasure-mediated mechanisms that seek to entice users into engaging, 3) disorientating interface elements that confuse users, block navigation, or inhibit their decision-making as a way to trap them into engaging, and 4) automated, low-friction designs that reduce user self-determination, encourage mindlessness, and lull users into engaging. These designs can occur in combination and may overlap in the psychological mechanisms that they leverage to achieve desired effects. Many can be delivered via recapture notifications in addition to the platform’s main interface. Here, we describe each of these four approaches in more depth.
@@ 4.1. Pressuring Users
Artificial Social Obligations. VLOPs translate users’ conscientiousness into extended use by presenting social obligations for the user to fulfill. This often takes the form of nudges to support or interact with other people, such as encouraging the user to congratulate a LinkedIn connection on a work anniversary or to give a virtual gift to another account on TikTok. In other instances, VLOPs encourage users to keep up with others’ content, nudge users to maintain interactions through artificial “streaks,” or notify users when someone has viewed their profile. They encourage users to create and post content by framing these activities as social obligations, saying things like “Your followers want to hear from you!” In all of these cases, the design of the platform works together with social anxiety and reciprocity bias to pressure users into engaging in interactions with other people that can only be carried out on the platform. Although notifications of these social obligations can be turned off on some VLOPs, doing so requires the user to take additional steps (e.g., LinkedIn requires four navigational steps to turn off connection-related notifications).
VLOPs also anthropomorphize their interfaces in ways that create social pressure to engage according to the platform’s interests. These anthropomorphic elements encourage the user to respond to the platform as if it were a person, creating pressure to treat the platform “well” (which the platform defines as extending engagement). For example, if a user attempts to delete their account on the Booking.com platform, a confirmshaming dialog asks, “Is this goodbye?” (see Figure 4), manufacturing a subtle guilt trip and conflating the act of preserving an account on the platform with the act of tending an interpersonal relationship.
@@ 4.2. Enticing Users
Separately, we found that VLOPs encourage reward- and pleasure-seeking via algorithms that prioritize engaging content. This incentivizes content creators to cater to impulsive, subconscious preferences that generate fast-brain clicks. This content is then organized within the interface to be maximally enticing, and when users step away from the platform, suspense-building notifications attempt to draw them back. Presenting potentially enticing content has the secondary benefit of allowing the platform to profile the user, learning their particular vulnerabilities by observing the types of clickbait—including those created by generative artificial intelligence—that are most effective in extending their engagement.
Reinforcement Schedules. VLOPs frequently employ interface elements that reward and affirm users for engaging with the platform or extending their time-on-task. In some instances, this involves offering rewards (such as free items, tokens, social rewards like “likes,” or pleasurable content) on intermittent or obscured reinforcement schedules, triggering dopaminergic pathways and increasing cravings to return to the app in hopes of better luck next time. In other instances, the platform reinforces prolonged engagement by rewarding it explicitly, for example, by rewarding daily logins, congratulating users on their activities within the app, or conferring badges or status as a function of engagement.
Curiosity Teasers. VLOPs also leverage gamification and suspense to increase usage by building users’ anticipation. For example, VLOPs hint at mystery content by showing just a glimpse of an item, displaying it partially off-screen and encouraging users to continue to scroll to satisfy their curiosity. In other cases, VLOPs promise information rewards via irresistible prompts like “See who’s viewed your profile” (Figure 6, left) or the promise of better deals if they continue exploring (Figure 6, middle). These teasers encourage the user to continue engaging to resolve a mystery that has been artificially constructed by the platform.
Figure 6. Curiosity teasers on different VLOPs. As users navigate through the platform, they are baited with teasers about who is looking at their profile (LinkedIn, left), encouraged to stay on the site to keep digging for the best deal (booking.com, middle), and promised mystery prizes as rewards for daily engagement (AliExpress, right).
@@
Tailoring Content to Base and Hyper-Personalized Pleasures. Additionally, VLOPs serve up guilty-pleasure recommendations to entice users with irresistible content. In these instances, the user does not search for or seek out this content; it is proactively delivered by the platform. Hypersexual and grotesque images extend users’ time on the platform by producing a feeling of being unable to look away. VLOPs bombard users with guilty-pleasure content in feeds, advertisements, popups, mainline content, and sidebar suggestions.
Figure 7. VLOPs attempt to extend engagement by appealing to base pleasures and enticing users with sexual, violent, or grotesque content.
@@
Social Quantification. We found that, in other instances, VLOPs entice users into extended engagement by appealing to their desire for social validation and connection. They provide features that enable affirmation from others, quantifying likes, followers, and other measures of responsiveness to the user’s content. This encourages the user to engage with the platform, first, by sharing content that others are likely to respond to, and second, by returning again and again to track the validation they may or may not be receiving from others. Like curiosity teasers, these social rewards are often embedded into push notifications (e.g., “You have two new followers”), enabling the platform to encourage engagement and redirect the user’s attention back to the platform, even when they are otherwise occupied.

### v2

@@ 2.3. Adolescent Development and Engagement-Prolonging Design
Understanding the manipulative designs that young people are exposed to is important, because there are theoretical reasons to expect that they may be disproportionately vulnerable to deception and pressure by platforms. Teens’ reduced impulse control and heightened sensitivity to rewards and social feedback make them particularly vulnerable to pressure and distraction (digital or otherwise). During adolescence, executive functions such as impulse inhibition, flexible problem-solving, and abstract reasoning develop at different rates between teens, with most of these higher-order thinking skills reaching maturity in the mid-20s. Therefore, teens are more susceptible to digital designs that exert manipulation through curiosity and novelty/risk-seeking, social pressure, and fear of missing out, with significant inter-individual variability  (of Sciences Engineering and Medicine, 2024).
@@
Teens’ accounts of their experiences with technology reflect frustration with attention-economy designs; more than half of teens say that they are online “almost constantly,” and approximately 40% of teens say they wish they spent less time using devices and social media (Center, 2018). As a result, many teens report attempting to cut back on these experiences. They delete apps that are too aggressive in demanding their attention (Landesman et al., 2024) and turn their account passwords over to their friends to lock them out of platforms that they find too addictive (Baumer et al., 2013). They add friction by burying apps in folders  (Böhmer and Krüger, 2013), setting do-not-disturb times  (Natarajan, 2024), and accessing platforms through browsers. But despite their makeshift strategies to resist the design tactics of the attention economy, teens still struggle to change their habits (Tran et al., 2019), suffer from device-based sleep disruption (Woods and Scott, 2016; Viner et al., 2019), and struggle with distraction from activities they value, like homework and time spent with friends and loved ones (Yardi and Bruckman, 2011). Parents of younger children report similar experiences and say that they frequently negotiate or argue with their child about transitioning away from media, especially media that uses engagement-prolonging designs like autoplay (Davis et al., 2019; Hiniker et al., 2016).
@@
In the midst of the ongoing debate regarding social media, technology, and youth mental health, there has been increasing focus on the role of engagement-prolonging design features as a mechanism linking platform use to reduced wellbeing (Abuse and, SAMHSA). Designs such as autoplay, endless scroll, and time pressure were described by the U.S. White House Interagency Kids Online Health and Safety Task Force as warranting modifications by platforms in order to reduce compulsive and excessive use of technology. Ongoing regulatory efforts have also targeted manipulative designs in products used by children and teens. For instance, the UK’s Online Safety Bill aims to make social media companies legally responsible for keeping children and young people safe online (Leg, 2023). And the UK’s Information Commissioner released the “Children’s Code Design Guidance” document, directing platforms to avoid exposing children and teens to engagement nudges or other EPDs (Information Commissioner’s Office (2025), ICO). These actions suggest widespread interest in regulating platforms’ ability to manipulate young people via EPDs. Our work seeks to contribute to these efforts by characterizing and quantifying the EPDs shown to teens on the world’s most commonly used platforms.
@@ 3. Method
We conducted an analysis of the web and app interfaces of all 17 VLOPs identified by the Digital Services Act (DSA) [^2] List of 17 VLOPs available on the European Commission’s website: https://ec.europa.eu/commission/presscorner/detail/en/IP_23_2413 , attending specifically to the presence of EPDs. We conducted this work in two phases (see Figure 2); in the first phase, we performed an inductive-deductive thematic analysis to identify and categorize the EPDs present in these interfaces, examining these designs in the context of the extant literature, and organizing them by the deeper strategies that underlie them. In the second phase, we conducted a structured content analysis of all VLOP interfaces to document all instances of the EPDs we identified in phase 1.
@@ 3.1. Sampling
We engaged extensively with each VLOP web interface ($N=17$) and app interface ($N=16$), in each case posing as a teenage user located in the European Union to enable researchers to engage more effectively with the design space with better imagination and empathy for the user experience (Medler and Magerko, 2010; Burns et al., 1994; Buchenau and Suri, 2000). We adopted this method because observing these platforms requires a significant amount of time, and we believe exposing teenagers to the manipulative design patterns presented on these platforms directly for a long period is problematic.
To observe websites, we used the Chrome browser in Guest Mode on a Windows 11 laptop. To observe mobile apps, we used a factory-reset smartphone running Android 9, equipped with an active SIM card. These devices were selected because they run the most widely used operating systems in Europe (StatCounter, 2024a, b) and therefore the operating systems that would be more commonly used by teenagers in Europe. We created two new Google accounts, which we used to access all VLOPs (we refer to these as the “primary account,” which was used to engage with all features, and the “secondary account,” which was used to enable the primary account to engage with social features that required a partner). The only exception to the above procedures was our analysis of the Apple App Store interface, which we observed using iOS devices and accounts due to the platform’s restrictions.
Following the method developed by Di Geronimo and colleagues (Di Geronimo et al., 2020), we collect screen-capture videos of each service on each modality that we could later examine in detail for evidence of EPDs. Both mobile and desktop interactions were recorded using Screen Recorder over Wi-Fi. Since platform companies may conduct A/B testing their products, the exact behaviors or interfaces we observed in a platform may not represent what all users would have encountered during the study period (the spring and summer of 2024).
@@
    Download App and Create an Account: We began by downloading the app and creating a new account with our teenager persona. The primary Google account was set up with the profile of a 17-year-old girl. Additionally, we registered a secondary account with the profile of a 17-year-old boy to interact with the primary account, as detailed in step four.
@@
While the screen recordings contained the visual captures of EPDs and interactions thereof, the audio served as complementary material, which allowed the coders to get a detailed impression of reviewers’ perception and judgement of a scene. We captured both screenshot and video records of the the user interfaces during these exploration and complied annotated images in a FigJam  (Fig, 2012) collective virtual workspace as structured data samples.
3.2. Content Analysis: Identifying the EPDs used by VLOPs
We then conducted a two-part analysis to identify the EPDs used by VLOPs. We first performed an inductive-deductive thematic analysis to develop a taxonomy of the EPDs used by VLOPs. We then conducted a structured content analysis to apply it comprehensively across platforms. Here, we describe these stages in detail.
@@ 3.2. Content Analysis: Identifying the EPDs used by VLOPs
    Inductive-Deductive Coding. We began our exploration with guidance from existing literature on attention-capture. In particular, we used the categories described by Radesky et al. (Radesky et al., 2022), Monge Roferello et al. (Monge Roffarello et al., 2023), and Flayelle et al. (Flayelle et al., 2023) as sensitizing concepts (Tracy, 2013) during our review of VLOP interfaces. We also allowed for emergent themes in EPDs and noted designs that did not fit neatly into these pre-existing taxonomies. Four authors independently reviewed the VLOP interfaces using this inductive-deductive approach and documented potential EPDs in a collective spreadsheet while interacting with VLOPs as described above. We also shared and organized exemplar screenshots in a collaborative Figma (Fig, 2012) board. Consistent with prior work (Di Geronimo et al., 2020; Gunawan et al., 2021), we made note of potential EPDs regardless of our perceptions of the designer’s intent, allowing us to include EPDs that may not reflect malicious intentions. The team met regularly over several months, and with each iteration brought new examples to discuss. During discussions, the team compared and contrasted designs, discussed possible groupings, assessed design examples against existing taxonomies, and refined possible category definitions. This process generated 38 inductive and 85 deductive EPD codes and associated examples.
@@
    Hierarchical Affinity Diagramming. The research team then conducted more formal affinity diagramming (Blandford et al., 2016) to restructure the codes and create an initial codebook. We structured our initial codebook into three levels: the psychological mechanisms through which the EPD design operates (high level), the recurring EPD design tactics (medium level), and the specific EPD design elements or interface components (low level). We added definitions and descriptions for each item at each level.
@@
    Inter-rater Reliability Testing. Two authors then independently reviewed two randomly selected VLOPs and applied the full set of codes across the interface. For each EPD, reviewers were required to note the code number of the mechanism and pattern, as well as specify the features exhibiting the EPDs (e.g., ”1(a) gifting option”). The coding criteria focused on the distribution of EPDs across screenshots or interactions extracted from the recordings, rather than the frequency of EPDs occurrences. In those cases where an aspect of the interface was deemed a potential EPDs but was not present within the existing codebook, reviewers proposed and added a new code as a comment with a visual example. Any confusion about specific cases or the taxonomy itself was recorded as a comment. Inter-rater reliability (IRR) was established and measured following Lombard et al.  (Lombard et al., 2002). The results of the first round yielded an IRR score below 0.6, indicating that the initial codebook (v1) fell below the minimum validity standard. Subsequently, the research team collaboratively compared the two internal coding results, discussed points of disagreement, unresolved questions, and confusion comments. This led to the development of codebook (v2), which included a refined definitions and names for each code, and the modification of codes. Codebook(v2) includes 71 EPD features in total. The same two researchers conducted a second round of internal testing with two different randomly chosen platforms, adhering to the same tasks with codebook (v2). The IRR score for this round was 0.715, exceeding the standard of 0.6 and indicating that the revised coding scheme and codebook (v2) were clear and interpretable enough to be the final version.
@@
    Assessing the Prevalence of EPDs across VLOPs. Finally, two authors independently coded all recordings from all VLOP interfaces for each of the EPDs we documented. They then compared their coding results and came up with the final data.
[…]
The image shows the final version of the EPD codebook, which includes categories and descriptions for the EPD framework.
Figure 3. Final version of the EPD codebook. More details can be found on the EPD website [^4] https://engagementprolongingdesign.github.io/EngagementProlongingDesign/.
4. Results Part 1: The EPDs Used by VLOPs and the Underlying Strategies Behind Them
We found that the EPDs that VLOPs employ fall into four broad categories, defined by the underlying manipulation strategy the design draws upon. These include: 1) artificial obligations manufactured by the platform that pressure users into engaging, 2) reward- or pleasure-mediated mechanisms that seek to entice users into engaging, 3) disorientating interface elements that confuse users, block navigation, or inhibit their decision-making as a way to trap them into engaging, and 4) automated, low-friction designs that reduce user self-determination, encourage mindlessness, and lull users into engaging. These designs can occur in combination and may overlap in the psychological mechanisms that they leverage to achieve desired effects. Many can be delivered via recapture notifications in addition to the platform’s main interface. Here, we describe each of these four approaches in more depth.
@@ 4.1. Pressuring Users
Artificial Social Obligations. VLOPs translate users’ conscientiousness into extended use by presenting social obligations for the user to fulfill. This often takes the form of nudges to support or interact with other people, such as encouraging the user to congratulate a LinkedIn connection on a work anniversary or to give a virtual gift to another account on TikTok. In other instances, VLOPs encourage users to keep up with others’ content, nudge users to maintain interactions through artificial “streaks,” or notify users when someone has viewed their profile. They encourage users to create and post content by framing these activities as social obligations, saying things like “Your followers want to hear from you!” In all of these cases, the design of the platform works together with social anxiety and reciprocity bias to pressure users into continuing to engage in interactions with other people that can only be carried out on the platform. Although notifications of these social obligations can be turned off on some VLOPs, doing so requires the user to take additional steps (e.g., LinkedIn requires four navigational steps to turn off connection-related notifications).
VLOPs also anthropomorphize their interfaces in ways that create social pressure to engage according to the platform’s interests. These anthropomorphic elements encourage the user to respond to the platform as if it were a person, creating pressure to treat the platform “well” (which the platform defines as extending engagement). For example, if a user attempts to delete their account on the Booking.com platform, a confirmshaming dialog asks, “Is this goodbye?” (see Figure  4), manufacturing a subtle guilt trip and conflating the act of preserving an account on the platform with the act of tending an interpersonal relationship.
@@ 4.2. Enticing Users
Separately, we found that VLOPs encourage reward- and pleasure-seeking via algorithms that prioritize engaging content. This incentivizes content creators to cater to impulsive, subconscious preferences that generate fast-brain clicks governed by ”System 1” thinking (Kahneman, 2011). This content is then organized within the interface to be maximally enticing, and when users step away from the platform, suspense-building notifications attempt to draw them back. Presenting potentially enticing content has the secondary benefit of allowing the platform to profile the user, learning their particular vulnerabilities by observing the types of clickbait—including those created by generative artificial intelligence—that are most effective in extending their engagement.
Reinforcement Schedules. VLOPs frequently employ interface elements that reward and affirm users for engaging with the platform or extending their time-on-task. In some instances, this involves offering rewards (such as free items, tokens, social rewards like “likes,” or pleasurable content) on intermittent or obscured reinforcement schedules, triggering dopaminergic pathways and increasing cravings to return to the app in hopes of better luck next time. In other instances, the platform reinforces prolonged engagement by rewarding it explicitly, for example, by rewarding daily logins, congratulating users on their activities within the app, or conferring badges or status as a function of engagement (Figure 6, left).
Curiosity Teasers. VLOPs also leverage gamification and suspense to increase usage by building users’ anticipation. For example, VLOPs hint at mystery content by showing just a glimpse of an item, displaying it partially off-screen and encouraging users to continue to scroll to satisfy their curiosity. In other cases, VLOPs promise information rewards via irresistible prompts like “See who’s viewed your profile” (Figure 6, middle) or the promise of better deals if they continue exploring (Figure 6, right). These teasers encourage the user to continue engaging to resolve a mystery that has been artificially constructed by the platform.
Figure 6. As users navigate through VLOPs, they are baited with promised mystery prizes as rewards for daily engagement (AliExpress, left), teasers about who is looking at their profile (LinkedIn, middle), encouraged to stay on the site to keep digging for the best deal (booking.com, right).
@@
Tailoring Content to Base and Hyper-Personalized Pleasures. Additionally, VLOPs serve up guilty-pleasure recommendations to entice users with irresistible content. In these instances, the user does not search for or seek out this content; it is proactively delivered by the platform. Hypersexual and grotesque images extend users’ time on the platform by producing a feeling of being unable to look away (see Figure 7). VLOPs bombard users with guilty-pleasure content in feeds, advertisements, popups, mainline content, and sidebar suggestions.
Figure 7. VLOPs attempt to extend engagement by appealing to base pleasures, enticing users with sexual, violent, or grotesque content. Left: AliExpress uses sexualized images to attract clicks and direct users to product links; Middle: Snapchat promotes violent user-challenge videos to capture attention; Right: TikTok leverages employs grotesque content to provoke reactions and encourage user interaction.
@@
Social Quantification. We found that, in other instances, VLOPs entice users into extended engagement by appealing to their desire for social validation and connection. They provide features that enable affirmation from others, quantifying likes, followers, and other measures of responsiveness to the user’s content (see Figure 8). This encourages the user to engage with the platform, first, by sharing content that others are likely to respond to, and second, by returning again and again to track the validation they may or may not be receiving from others. Like curiosity teasers, these social rewards are often embedded into push notifications (e.g., “You have two new followers”), enabling the platform to encourage engagement and redirect the user’s attention back to the platform, even when they are otherwise occupied.